# Taming Rectified Flow for Inversion and Editing

### v1

@@ Abstract
Rectified-flow-based diffusion transformers, such as FLUX and OpenSora, have demonstrated exceptional performance in the field of image and video generation. Despite their robust generative capabilities, these models often suffer from inaccurate inversion, which could further limit their effectiveness in downstream tasks such as image and video editing. To address this issue, we propose RF-Solver, a novel training-free sampler that enhances inversion precision by reducing errors in the process of solving rectified flow ODEs. Specifically, we derive the exact formulation of the rectified flow ODE and perform a high-order Taylor expansion to estimate its nonlinear components, significantly decreasing the approximation error at each timestep. Building upon RF-Solver, we further design RF-Edit, which comprises specialized sub-modules for image and video editing. By sharing self-attention layer features during the editing process, RF-Edit effectively preserves the structural information of the source image or video while achieving high-quality editing results. Our approach is compatible with any pre-trained rectified-flow-based models for image and video tasks, requiring no additional training or optimization. Extensive experiments on text-to-image generation, image & video inversion, and image & video editing demonstrate the robust performance and adaptability of our methods. Code is available at this URL.
@@
Figure 1: We propose RF-Solver to solve the rectified flow ODE with less error, thus enhancing both sampling quality and inversion-reconstruction accuracy for rectified-flow-based generative models. Furthermore, we propose RF-Edit to leverage the RF-Solver for image and video editing tasks. Our methods achieve impressive performance on various tasks, including text-to-image generation, image & video inversion, and image & video editing.
@@ 1 Introduction
Recently, rectified-flow-based generation methods [31] have demonstrated remarkable performance in generating high-quality images and videos. Compared with traditional approaches such as Stable Diffusion [19, 42], their impressive capabilities stem primarily from two aspects. First, rectified flow utilizes a straightforward Ordinary Differential Equation (ODE), which constructs a continuous straight-line motion system to produce the desired data distribution. This simplified training and inference paradigm enables the model to more effectively learn the underlying distribution of real image data. Second, most of the latest rectified flow methods employ the Diffusion Transformer (DiT) [38, 59] architecture as the backbone, which has shown superior performance compared to traditional U-Net [43] architectures. As a result, image generation models like FLUX [1] and video generation models like OpenSora [2], both based on rectified flow and DiT, have respectively become one of the state-of-the-art (SOTA) models in Text-to-Image (T2I) and Text-to-Video (T2V) generation.
Apart from fundamental T2I and T2V tasks, other downstream tasks such as reconstruction [46, 35, 16] and editing [17] have been attracting growing interest. These tasks typically depend on performing inversion on the source image/video. The inversion process yields the corresponding representation in the noise space, which is followed by denoising with various conditions provided by users. Given the robust generative capabilities of rectified-flow-based models, they are expected to exhibit superior performance on these downstream tasks. However, their performance remains unsatisfactory compared with traditional methods (based on DDPM and UNets), and research in this area still lags behind.
Delving into this problem, we identify that the primary challenge lies in the significant errors during the inversion process of the rectified flow, which fail to accurately reconstruct the original images or videos (Task 1 and Task 2 in Fig. 1). This limitation further constrains its performance on other downstream tasks such as editing. Although [45] attempted to address this challenge through dynamic optimal control, its application is restricted to simple image editing scenarios such as stylization and facial editing, which typically involve simple content and uniform backgrounds. In contrast, both the image contents and editing requirements (such as addition, replacement, and global editing) are considerably more complex in the real world. Moreover, video editing requires highly consistent temporal modeling, presenting even greater challenges for editing algorithms. To the best of our knowledge, there are still no existing methods that effectively tackle these issues.
Instead of focusing on designing a specific inversion method, we aim to address the above problem from a more general and fundamental perspective: the sampler. This is because the essence of the inversion and generation process for rectified flow is to employ a sampler that estimates the solution of rectified flow ODE. Consequently, the primary source of inaccurate inversion lies in the approximation error in the solution, which accumulates at each timestep. Intuitively, if the ODE is solved more accurately, the accuracy of inversion can be enhanced subsequently.
@@
Based on this insight, we propose the RF-Solver. Specifically, we note that the ODE formulation for rectified flow can be solved directly using the variation of constants method, yielding an exact formulation of the solutions. For the nonlinear component of this solution (i.e., the integral of the neural network), we utilize Taylor expansion for estimation. By employing higher-order Taylor expansion, the ODE can be solved with reduced error, thereby improving the performance of rectified flow models. RF-Solver is a generic sampler that can be seamlessly integrated into any rectified flow model without the need for training or optimization. Experimental results demonstrate that RF-Solver not only significantly enhances the accuracy of inversion and reconstruction, but also improves performance on fundamental tasks such as T2I generation.
@@
Building upon this, we propose RF-Edit to apply RF-Solver in editing tasks. Real-world image and video editing require the model to make precise modifications to a source image/video while maintaining its overall structure unchanged. This makes editing a more challenging task compared to reconstruction. In this scenario, it is inadequate to solely rely on the inverted noises as prior knowledge for editing, which could lead to edited results being excessively influenced by the target prompt, resulting in a completely different output compared with source image/video. Addressing this problem, we store the $V$ (value) feature in the self-attention layers at several timesteps during inversion. In the process of denoising, these features are used to replace the corresponding features. Practically, we design two specific sub-modules for RF-Edit, respectively leveraging FLUX [1] and OpenSora [2] as the backbone for image and video editing. With the effective design of RF-Edit, it demonstrates superior performance in both image and video domains, outperforming various SOTA methods.
@@
  • We propose RF-Solver, a training-free sampler that significantly reduces errors in the inversion and reconstruction processes of the rectified-flow model.
  • We present RF-Edit, which leverages RF-Solver for image and video editing, effectively preserving the structural integrity of the source image/video while achieving high-quality results.
  • Extensive experiments on images and videos demonstrate the efficacy of our methods, showcasing superior performance in both inversion and high-quality editing compared to various existing baselines.
@@ 2.1 Inversion
Inversion maps the real visual data, i.e. image and video, to a representation in the noise space [14, 50, 35, 44, 34, 23], which is the reverse process of generation. This representation captures the essential features and structures of the original data while allowing for flexibility in manipulation for various editing applications. Numerous previous inversion approaches have been elaborately designed for diffusion models to achieve remarkable performance. The representative method, DDIM inversion [46, 47], adds predicted noise recursively at each forward step, outputting the final state as structured noise. To mitigate the discretization error in DDIM inversion, some efforts have been explored from the perspective of optimizing null prompt embeddings [35] and latent variables [44]. Negative prompt inversion [34] accelerates the inversion process at the cost of fidelity. Despite the success of inversion in diffusion models, the exploration of inversion in SOTA rectified flow models like FLUX and OpenSora is limited. RF-prior [57] uses the score distillation to invert the image while it requires a number of optimizing steps. More recently, [45] introduces an additional vector field that is conditioned on the source image to improve the inversion. However, its performance improvement mainly stems from the information leakage of the source image during the inversion and reconstruction process. The error from the original vector field of rectified flow still persists, which would limit the performance of such method on various downstream tasks.
@@ 2.2 Image and Video Editing
Training-free methods for image and video editing [21, 48] have gained increasing popularity due to their efficiency and effectiveness. These methods usually invert the source image/video into Gaussian noises and then denoises it conditioned on the target prompt. Existing image editing methods focus on prompt refinement [41, 52], attention-sharing mechanism [17, 37, 7, 49], mask guidance [5, 11, 20, 28], and noise initialization [6, 58]. Video editing introduces additional complexities due to the need for maintaining temporal consistency, making it a more challenging task. Existing video editing methods focus on attention injection [39, 53, 30], motion guidance [10, 15, 56, 51], latent manipulation [61, 55, 25, 9], and canonical representation [8, 27, 36, 26]. To date, the editing performance of rectified-flow-based diffusion transformers has remained largely under-explored. Although [45] employs FLUX for image editing, its performance is limited to simple tasks such as stylization and face editing, and it struggles to effectively maintain the structural information of source images. Moreover, currently there is no research exploring the video editing capabilities of rectified-flow-based models.
@@ 3 Method
In this section, we present our methods in detail. First, we introduce the proposed RF-Solver, which significantly enhances the precision of inversion and reconstruction. Subsequently, we introduce RF-Edit, an extension of RF-Solver designed to enable high-quality image and video editing.
@@ 3.1 Preliminaries
Rectified Flow [32] facilitates the transport between the Gaussian Noises distributions $\pi_{0}$ and real data distribution $\pi_{1}$ along a straight path. This is achieved by learning a forward-simulating system given by $\frac{d\boldsymbol{Z}_{t}}{dt}=v(\boldsymbol{Z}_{t},t),t\in[0,1]$ which maps $\boldsymbol{Z}_{0}\in\pi_{0}$ to $\boldsymbol{Z}_{1}\in\pi_{1}$. In practice, the velocity field $v$ is parameterized by a neural network $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t},t)$. During training, given $\boldsymbol{X}_{0}\in\pi_{0}$ and $\boldsymbol{X}_{1}\in\pi_{1}$, the forward process of rectified flow is implemented through a simple linear combination:
$\boldsymbol{X}_{t}=t\boldsymbol{X}_{1}+(1-t)\boldsymbol{X}_{0},t\in[0,1].$ (1)
The rectified flow ODE can be derived by differentiating Eq. 1 with respect to $t$, as follows:
$\frac{d\boldsymbol{X}_{t}}{dt}=\boldsymbol{X}_{1}-\boldsymbol{X}_{0},t\in[0,1].$ (2)
Here, $\boldsymbol{X}_{1}-\boldsymbol{X}_{0}$ serves as the ground truth, and the network is optimized through mean square error:
@@
During the sampling process, the ODE is discretized and solved using Euler method. Specifically, the rectified flow model starts with Gaussian noises $\boldsymbol{Z}_{t_{N}}\in\mathcal{N}(0,\boldsymbol{I})$. Given a series of discrete $N$ timesteps $t=\{t_{N},...,t_{0}\}$, the model iteratively predicts $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}}),i\in\{N,\cdots,1\}$ and then takes a step forward until generating the images $\boldsymbol{Z}_{t_{0}}$, with the following recurrence relation:
@@
The definition and training process of Rectified Flow is designed to establish a nearly linear transition trajectory between two distributions, enabling efficient generation with significantly fewer steps compared to DDPMs [19].
@@ 3.2 RF-Solver
Figure 2: Analysis of the Inversion-Reconstruction Process. Inversion begins with the source image latent $\widetilde{\boldsymbol{{Z}}}_{t_{0}}$ as the input and progressively add noise, obtaining $\widetilde{\boldsymbol{{Z}}}_{t_{N}}\in\mathcal{N}(0,\boldsymbol{I})$. $\widetilde{\boldsymbol{Z}}_{t_{N}}$ is then denoised for $N$ timesteps to obtain the reconstruction $\boldsymbol{Z}_{t_{0}}$. During this process, we store the latent $\boldsymbol{\widetilde{Z}}_{t_{i}}$ and $\boldsymbol{Z}_{t_{i}}$ at each time step and calculate the mean squared error between them.
Figure 3: RF-Edit Pipelines. The RF-Edit framework comprises two sub-modules, respectively applied to FLUX [1] for image editing and OpenSora [2] for video editing. For image editing, we share the feature of $\widetilde{\mathcal{V}}$ within the single block self-attention of the FLUX [1] backbone. For video editing, we share the feature of $\widetilde{\mathcal{V}}$ in the spatial attention of the OpenSora [2] backbone.
The vanilla rectified flow (RF) sampler demonstrates strong performance in image and video generation. However, when applied to inversion and reconstruction tasks, we observe significant error accumulation at each timestep. This results in reconstructions that diverge notably from the original image (see Fig. 2), further limiting the performance of RF-based models in various downstream tasks, such as image and video editing. Delving into this problem, we notice that the inversion and reconstruction processes in rectified flow rely on estimating an approximate solution of the rectified flow ODE at each timestep (see Eq. 4). Obtaining more precise solutions for the ODE would effectively mitigate these errors, leading to improved reconstruction quality. Based on this analysis, we start by carefully examining the differential form of the Rectified flow: $\frac{d\boldsymbol{Z}_{t}}{dt}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t},t)$. This ODE is discretized in the sampling process. Given the initial value $\boldsymbol{Z}_{t_{i}}$, the ODE can be exactly formulated using the variant of constant method:
@@
Through the above process, the network prediction term and its higher-order derivatives are separated from the integral. Then we notice that the remaining portion in the integral can be computed analytically:
@@
  $\boldsymbol{v}_{t_{i}}^{(1)}\leftarrow({\boldsymbol{\hat{v}}_{t_{i}}-\boldsymbol{\hat{v}}_{t_{i}+\Delta t_{i}}})/{\Delta t_{i}}$triangleright Calculating the Derivatives
@@
When $n=1$, the formula reduces to the standard rectified flow (Eq. 4). In our experiments, we find that setting $n=2$ effectively mitigates the errors, yielding the following formula:
Note that $\boldsymbol{v}_{\theta}^{(1)}$ is the first-order derivative of the network prediction term $\boldsymbol{v}_{\theta}$, which cannot be analytically derived due to the complex architecture of the neural network. To estimate this term, we first obtain the network prediction $\hat{\boldsymbol{v}}_{t_{i}}$ at the timestep $t_{i}$, i.e., $\hat{\boldsymbol{v}}_{t_{i}}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$. Then we step forward a small timestep $\Delta t=\frac{1}{2}(t_{i-1}-t_{i})$, and update the latents to obtain $\boldsymbol{Z}_{{t_{i}}+\Delta t}=\boldsymbol{Z}_{t_{i}}+\Delta t\cdot\hat{\boldsymbol{v}}_{t_{i}}$. Next, we calculate an additional prediction of the network at the timestep $t_{i}+\Delta t$, i.e., $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}+\Delta t},{t_{i}+\Delta t})$. With $\hat{\boldsymbol{v}}_{t_{i}}$ and $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}$, the first-order derivative of ${\boldsymbol{v}}_{\theta}$ at the timestep $t_{i}$ can be estimated as:
$\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i})=\frac{\hat{\boldsymbol{v}}_{t_{i}+\Delta t}-\hat{\boldsymbol{v}}_{t_{i}}}{\Delta t}.$ (11)
Substituting Eq. 11 into Sec. 3.2 results in the practical implementation of the RF-Solver algorithm. The complete sampling process for RF-Solver is presented in Algorithm 1.
Besides sampling, inversion seeks to map data back into noise, which reverses the sampling process. Following previous methods for DDIM inversion [46, 13], the ODE process can be directly reversed in the limit of small steps. Based on this assumption, the inversion process of RF-Solver (Sec. 3.2) can be directly transformed as:
@@
where $\widetilde{\boldsymbol{Z}}_{t_{i}}$ and $\widetilde{\boldsymbol{Z}}_{t_{i+1}}$ denotes the latents during inversion. Through this high order expansion, the error of the ODE solution in each timestep is reduced from $\mathcal{O}\big{(}(h_{i})^{2}\big{)}$ to $\mathcal{O}\big{(}(h_{i})^{3}\big{)}$, leading to improved performance, particularly in inversion and reconstruction (see Fig. 2). Beyond inversion and reconstruction, RF-Solver can also be applied to any RF-based model (such as FLUX [1] and OpenSora [2]) for other tasks such as sampling and editing, enhancing performance without requiring additional training.

### v2

@@ Abstract
Rectified-flow-based diffusion transformers like FLUX and OpenSora have demonstrated outstanding performance in the field of image and video generation. Despite their robust generative capabilities, these models often struggle with inversion inaccuracies, which could further limit their effectiveness in downstream tasks such as image and video editing. To address this issue, we propose RF-Solver, a novel training-free sampler that effectively enhances inversion precision by mitigating the errors in the ODE-solving process of rectified flow. Specifically, we derive the exact formulation of the rectified flow ODE and apply the high-order Taylor expansion to estimate its nonlinear components, significantly enhancing the precision of ODE solutions at each timestep. Building upon RF-Solver, we further propose RF-Edit, a general feature-sharing-based framework for image and video editing. By incorporating self-attention features from the inversion process into the editing process, RF-Edit effectively preserves the structural information of the source image or video while achieving high-quality editing results. Our approach is compatible with any pre-trained rectified-flow-based models for image and video tasks, requiring no additional training or optimization. Extensive experiments across generation, inversion, and editing tasks in both image and video modalities demonstrate the superiority and versatility of our method. The source code is available at this URL.
@@
Figure 1: RF-Solver for downstream tasks in image and video. We propose RF-Solver to solve the rectified flow ODE with reduced error, thereby enhancing both sampling quality and inversion-reconstruction accuracy for rectified-flow-based generative models [2, 3]. Furthermore, we propose RF-Edit, which utilizes the RF-Solver for image and video editing tasks. Our methods demonstrate impressive performance across generation, inversion, and editing tasks in both image and video modalities.
@@ 1 Introduction
Recent advancements of generation methods based on Rectified Flow (RF) [42, 35, 67] have demonstrated exceptional performance in synthesizing high-quality images and videos. Different from traditional approaches represented by Stable Diffusion [22, 56], these methods leverage the Diffusion Transformer [51, 78, 73, 63] architecture and implement a straight-line motion system to produce the desired data distribution. With these effective designs, FLUX [2] and OpenSora [3] have respectively emerged as one of the state-of-the-art (SOTA) methods in the field of Text-to-Image (T2I) and Text-to-Video (T2V) generation.
Despite the remarkable success in the fundamental T2I and T2V generation tasks, few studies have explored the performance of RF-based models on various downstream tasks such as inversion-reconstruction [60, 47, 17, 71] and editing [20, 45]. When directly applying the vanilla RF for inversion, we observe that it fails to faithfully reconstruct the image or video from the source. Examples are shown in Fig. 1 Task 1 and Task 2 (the third row). For image inversion, the positions of objects (e.g. the cake) and the appearance of individuals (e.g. the child) in the reconstructed image significantly diverge from the source image. The performance of video inversion is even worse, with noticeable distortions present in the reconstructed video. The inaccuracies of inversion and reconstruction would severely constrain the performance of RF models on other inversion-based downstream tasks such as image editing [20, 64, 48, 13, 29] and video editing [41, 33, 15, 59].
In this work, we investigate the aforementioned problem by delving into the inversion and reconstruction process of the RF. Specifically, we track the latent at each intermediate timestep during inversion and reconstruction, calculating the Mean Square Error (MSE) between them at corresponding timesteps. We observe that significant errors are introduced at each timestep throughout the whole reconstruction process, and their accumulation ultimately results in a considerably deviated output (the red curve in Fig. 2). Based on the definition and inference process of RF [40, 42], we identify that these errors stem from the Ordinary Differential Equation (ODE) solving process [79, 66, 23, 52]. Specifically, the essence of the inversion and generation process for RF is to derive the solution of RF ODE [42]. Since this ODE includes terms involving complex neural networks, the solution can only be coarsely approximated by a sampler. However, the experiment in Fig. 2 indicates that the sampler adopted in existing models [2, 3] lacks sufficient precision for the inversion task, causing notable errors to accumulate at each timestep, finally leading to unsatisfactory reconstruction results.
Figure 2: Analysis of the inversion-reconstruction process. Inversion takes the source image latent $\widetilde{\boldsymbol{{Z}}}_{t_{0}}$ as the input and progressively add noise for $N$ timesteps, obtaining $\widetilde{\boldsymbol{{Z}}}_{t_{N}}\in\mathcal{N}(0,\boldsymbol{I})$. $\widetilde{\boldsymbol{Z}}_{t_{N}}$ is then denoised for $N$ timesteps to obtain the reconstruction $\boldsymbol{Z}_{t_{0}}$. During this process, we store the latent $\boldsymbol{\widetilde{Z}}_{t_{i}}$ and $\boldsymbol{Z}_{t_{i}}$ at each timestep respectively in inversion and denoising processes. Then we calculate the Mean Squared Error (MSE) between them. The red curve represents the vanilla Rectified Flow inversion and the green curve represents RF-Solver inversion.
@@
Based on the analysis, we aim to improve inversion accuracy by introducing a more effective sampler, which is more general and fundamental than designing a specific inversion method. To this end, we propose RF-Solver. Specifically, we note that the exact formulation of the RF ODE solution can be directly derived using the variation of constants method. For the nonlinear component of this solution (i.e., the integral of the neural network), we utilize Taylor expansion for estimation. By employing higher-order Taylor expansion, the ODE can be solved with reduced error, thereby enhancing the performance of RF models. RF-Solver is a generic sampler that can be seamlessly integrated into any rectified flow model without additional training or optimization. Experimental results demonstrate that RF-Solver not only significantly enhances the accuracy of inversion and reconstruction (the green curve in Fig. 2), but also improves performance on fundamental tasks such as T2I generation.
@@
Building upon this, we propose RF-Edit to leverage RF-Solver in editing tasks. Real-world image and video editing require the model to make precise modifications to a source image/video while maintaining its overall structure unchanged, presenting greater challenges than reconstruction. In this scenario, it is inadequate to solely rely on the inverted noises as prior knowledge for editing, which could lead to edited results being excessively influenced by the target prompt, diverging significantly from the original source [20, 64]. Addressing this problem, RF-Edit stores the $\mathcal{V}$ (value) feature in the self-attention layers at several timesteps during inversion. These features are used to replace the corresponding features in the denoising process. Practically, we design two specific sub-modules for RF-Edit, respectively leveraging the DiT structure of FLUX [2] and OpenSora [3] as the backbones for image and video editing. With the effective design of RF-Edit, it demonstrates superior performance in both image and video domains, outperforming various SOTA methods.
@@
  • We propose RF-Solver, a training-free sampler that significantly reduces errors in the inversion and reconstruction processes of rectified-flow models.
  • We present RF-Edit, which leverages RF-Solver for image and video editing. RF-Edit effectively preserves the structural integrity of the source image/video while delivering high-quality results.
  • Extensive experiments across a range of tasks demonstrate the efficacy and versatility of our method.
@@ 2.1 Inversion
Inversion maps the real visual data, i.e. image and video, to representations in noise space, which is the reverse process of generation. The representative method, DDIM inversion [60, 61], adds predicted noise recursively at each forward step. Many efforts [14, 65, 47, 57, 46, 27, 46, 44, 19] have been made to mitigate the discretization error in DDIM inversion. Despite the effectiveness of inversion in diffusion models, the exploration of inversion in SOTA rectified flow models like FLUX and OpenSora is limited. RF-prior [76] uses the score distillation to invert the image while it requires many optimizing steps. More recently, [58] introduces an additional vector field conditioned on the source image to improve the inversion. However, the error from the original vector field of rectified flow still persists, which would limit the performance of such method on various downstream tasks. In contrast, we aim to directly mitigate the error from the original vector field in this work.
@@ 2.2 Image and Video Editing
Training-free methods for image and video editing [25, 62] have gained increasing popularity for their efficiency and effectiveness. Existing image editing methods focus on prompt refinement [55, 69], attention-sharing mechanism [20, 50, 18, 64], mask guidance [4, 10, 24, 37, 19], and noise initialization [5, 77]. Video editing introduces additional complexities in maintaining temporal consistency, making it a more challenging task. Existing video editing methods focus on attention injection [53, 70, 41], motion guidance [9, 16, 75, 68], latent manipulation [81, 74, 30, 8], and canonical representation [7, 36, 49, 31]. To date, the editing performance of RF-based diffusion transformers has remained largely under-explored. Although [58] employs FLUX [2] for image editing, its performance is limited to simple tasks such as stylization and face editing while often failing to effectively maintain the structural information of source images. Moreover, currently there is no research exploring the video editing capabilities of RF-based models.
@@ 3 Method
In this section, we present our method in detail. First, we introduce RF-Solver, which significantly enhances the precision of inversion and reconstruction. Subsequently, we present RF-Edit, an extension of RF-Solver designed to enable high-quality image and video editing.
@@ 3.1 Preliminaries
Rectified Flow (RF) [43] facilitates the transition between the real data distribution $\pi_{0}$ and Gaussian Noises distribution $\pi_{1}$ along a straight path. This is achieved by learning a forward-simulating system defined by the ODE: $d\boldsymbol{Z}_{t}=v(\boldsymbol{Z}_{t},t)dt,t\in[0,1]$, which maps $\boldsymbol{Z}_{1}\in\pi_{1}$ to $\boldsymbol{Z}_{0}\in\pi_{0}$.
In practice, the velocity field $v$ is parameterized by a neural network $\boldsymbol{v}_{\theta}$. During training, given empirical observations of two distributions $\boldsymbol{X}_{0}\sim\pi_{0}$, $\boldsymbol{X}_{1}\sim\pi_{1}$ and $t\in[0,1]$, the forward process (i.e. adding noise) of rectified flow is defined by a simple linear combination: $\boldsymbol{X}_{t}=t\boldsymbol{X}_{1}+(1-t)\boldsymbol{X}_{0}$. The differential form of the equation is given by: $d\boldsymbol{X}_{t}=(\boldsymbol{X}_{1}-\boldsymbol{X}_{0})dt$. Consequently, the training process optimizes the network by solving the least squares regression problem, which fits the $\boldsymbol{v}_{\theta}$ with $(\boldsymbol{X}_{1}-\boldsymbol{X}_{0})$:
@@
In the sampling process, the ODE is discretized and solved using the Euler method. Specifically, the rectified flow model starts with a Gaussian noise sample $\boldsymbol{Z}_{t_{N}}\in\mathcal{N}(0,\boldsymbol{I})$. Given a series of $N$ discrete timesteps $t=\{t_{N},...,t_{0}\}$, the model iteratively predicts $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$ for $i\in\{N,\cdots,1\}$ and then takes a step forward until generating the images $\boldsymbol{Z}_{t_{0}}$, with the following recurrence relation:
@@
The RF model can generate high-quality images in much fewer timesteps compared to DDPM [22], owing to the nearly linear transition trajectory established during training. With these effective designs, RF model illustrates great potential in the field of T2I and T2V generation [2, 3].
@@ 3.2 RF-Solver
The vanilla RF sampler demonstrates strong performance in image and video generation. However, when applied to inversion and reconstruction tasks, we observe significant error accumulation at each timestep. This results in reconstructions that diverge notably from the original image (see Fig. 2). This severely limits the performance of RF models in various inversion-based downstream tasks [20, 66]. Delving into this problem, we identify that the errors stem from the process of estimating the approximate solution for the rectified flow ODE [67, 40], which is formulated by Eq. 2 in existing methods [2, 3]. Consequently, obtaining more precise solutions for the ODE would effectively mitigate these errors, leading to improved reconstruction quality.
Based on this analysis, we start by carefully examining the differential form of the Rectified flow: ${d\boldsymbol{Z}_{t}}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t},t)dt$. This ODE is discretized in the sampling process. Given the initial value $\boldsymbol{Z}_{t_{i}}$, the ODE can be exactly formulated using the variant of constant method:
@@
Through the above process, the network prediction term and its higher-order derivatives are separated from the integral. Then we notice that the remaining component in the integral can be computed analytically:
@@
When $n=1$, the formula reduces to the standard rectified flow (i.e., Eq. 2). In our experiments, we find that setting $n=2$ effectively mitigates the errors, yielding:
Figure 3: RF-Edit pipelines for image editing and video editing. We design two sub-modules for applying RF-Edit to (a). Image editing with FLUX [2] and (b). Video editing with OpenSora [3]. Note that for FLUX, there are multiple Double Blocks, followed by multiple Single Blocks. For OpenSora, there are multiple OpenSora DiT blocks. For simplicity, only one block of each type is depicted in the figure.
Note that $\boldsymbol{v}_{\theta}^{(1)}$ in Sec. 3.2 is the first-order derivative of the network prediction term $\boldsymbol{v}_{\theta}$, which cannot be analytically derived due to the complex architecture of the neural network. To estimate this term, we first obtain the network prediction $\hat{\boldsymbol{v}}_{t_{i}}$ at the timestep $t_{i}$, i.e., $\hat{\boldsymbol{v}}_{t_{i}}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$. Then we step forward a small timestep $\Delta t=\frac{1}{2}(t_{i-1}-t_{i})$, and update the latents to obtain $\boldsymbol{Z}_{{t_{i}}+\Delta t}=\boldsymbol{Z}_{t_{i}}+\Delta t\cdot\hat{\boldsymbol{v}}_{t_{i}}$.
Subsequently, we calculate an additional prediction of the network at the timestep $t_{i}+\Delta t$, i.e., $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}=\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}+\Delta t},{t_{i}+\Delta t})$. With $\hat{\boldsymbol{v}}_{t_{i}}$ and $\hat{\boldsymbol{v}}_{t_{i}+\Delta t}$, the first-order derivative of ${\boldsymbol{v}}_{\theta}$ at the timestep $t_{i}$ can be estimated as: $\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i})=\frac{\hat{\boldsymbol{v}}_{t_{i}+\Delta t}-\hat{\boldsymbol{v}}_{t_{i}}}{\Delta t}$. Substituting this formulation into Sec. 3.2 results in the practical implementation of the RF-Solver algorithm. The complete sampling process for RF-Solver is presented in Algorithm 1.
@@
  $\boldsymbol{v}_{t_{i}}^{(1)}\leftarrow({\boldsymbol{\hat{v}}_{t_{i}+\Delta t_{i}}-\boldsymbol{\hat{v}}_{t_{i}}})/{\Delta t_{i}}$triangleright Calculating the Derivatives
@@
Obtaining the sampling form of RF-Solver, we further derive its inversion form. Inversion maps data back into noise, which reverses the sampling process. Following previous methods for DDIM inversion [60, 12], the ODE process can be directly reversed in the limit of small steps. Based on this assumption, the inversion process of RF-Solver (Sec. 3.2) can be directly derived as:
@@
where $\widetilde{\boldsymbol{Z}}_{t_{i}}$ and $\widetilde{\boldsymbol{Z}}_{t_{i+1}}$ denotes the latents during inversion. Through the high order expansion, the error of the ODE solution in each timestep is reduced from $\mathcal{O}\big{(}(h_{i})^{2}\big{)}$ to $\mathcal{O}\big{(}(h_{i})^{3}\big{)}$, leading to improved performance, particularly in inversion and reconstruction (see Fig. 2). Beyond inversion and reconstruction, RF-Solver can also be applied to any RF-based model (such as FLUX [2] and OpenSora [3]) for other tasks such as sampling and editing, enhancing performance without requiring additional training.